# FSFM: A Generalizable Face Security Foundation Model via Self-Supervised Facial Representation Learning

### v1

@@ 4.3 Face anti-spoofing
Comparison with Base Models Our FSFM achieves the best average performance among base models, with several key observations: 1) Simple finetuning on ViT-B Sup(IN) enhances FAS generalization, as noted in [54, 134]. 2) General SSL methods, including MIM-based MAE [45] and ID-based DINO [11], perform worse than ViT-B Sup(IN), with MAE showing notable declines, echoing the claims in [120]. 3) With large-scale facial data for SSL, MCF outperforms MAE and DINO for FAS. However, similar to DfD, it still falls short of the ViT-B Sup(IN) baseline in average metrics, once again underscoring the gap between normal face analysis and security tasks. 4) Overall, our FSFM improves the generalizability of ViT for cross-domain FAS, effectively modeling credible features of live (real) faces.
Comparison with SOTA Specialized Methods Our FSFM ViT-B exhibits superior performance over domain generalization (DG) FAS methods using various CNNs. Notably, this comparison adheres to the baseline [54]: although supplementary data (CelebA-Spoof [124]) is included for fine-tuning ViT-B, no auxiliary supervision or DG techniques, e.g., depth map (DM) or adversarial learning (AL), are employed—only a standard cross-entropy loss. Additionally, we follow other ViT-based FAS methods [134, 67, 106] without including CelebA-Spoof, with comparisons provided in the supplementary material.
@@ 4.5 Ablation studies
We conduct extensive ablations to evaluate the effectiveness of each component. In this subsection, unless otherwise stated, we pretrain the ViT-B/16 network on the FF++_o dataset, which includes 0.11M real face images extracted from 720 training and 140 validation videos in the FF++ (c23) Youtube subset [84]. We present the average generalization performance on DfD and FAS in Tab. 4.
@@
Effect of Key Designs 1) Blocks of & A lightweight 2-layer Rep decoder in both the online and target branches (2&2) achieves a better balance between complexity and generalization, outperforming 1&1 and 3&3. Omitting the Rep decoder (0&0) or adding it only to the target branch (2&0) degrades performance, supporting the importance of bridging feature distribution in a disentangled space. 2) Data Augmentation FSFM performs best without data augmentation (none&none). Unlike many SSL methods, applying simple augmentation in the MIM network (crop+flip&none) or target view (none&crop+flip) reduces performance, indicating that CRFR-P masking introduces sufficient spatial regularization and that preserving target representations from original faces aids robustness. 3) Loss for ID Network Pretraining with asymmetric negative cosine similarity (NCS) i.e. Eq. 4 consistently outperforms the widely used InfoNCE, as the latter pushes different real faces (negative samples) apart that may not be suitable for face security tasks. Besides, our SimSiam-like [15] NCS proves more effective than BYOL-like [36] MSE.
@@ 4.6 Qualitative analysis
To illustrate the superiority of the facial features learned by FSFM, we visualize GradCAM [86] activation maps for detecting forgery and spoofing, compared to Sup(IN) and MAE baselines. 1) DfD In Fig. 4 (a), FSFM heatmaps more accurately identify forgery-relevant artifacts in FF++ [84], e.g., mouth-modified F2F and NT. 2) FAS In Fig. 4 (b), FSFM predominantly captures spoof-specific clues in the MCIO cross-domain testings: high-frequency paper textures (M-Paper), photo cut edges (C-Photo), inconsistent reflections between facial regions (I-Print), and moiré patterns in video (O-Replay). These visualizations show that FSFM effectively responds to anomalies diverging from the suggested 3C in real/live faces, shedding light on improving the generalization of face security tasks.
@@ 5 Conclusion
In this work, we propose a self-supervised pretraining framework to learn fundamental and general facial representations of real faces, FSFM. We leverage the synergy between masked image modeling and instance discrimination to empower both local and global perception, promoting intra-region consistency, inter-region coherency, and local-to-global correspondence. We show that our FSFM transfers better than previous pretrained models on several face security tasks, including cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion face forgery detection. Notably, as a vanilla ViT, our model surpasses SOTA task-specific methods on generalization performance. Finally, we hope that our work can contribute to the face security-related research.
@@ A Overview
This supplementary material provides additional insights, details, and results supporting our FSFM framework, structured as follows:
Facial Masking Strategies in MIM (Sec. B): We delve into the impact of different facial masking strategies on naive MAE, including quantitative and qualitative analyses of attention differences.
Instance Discrimination in FSFM (Sec. C): We highlight the connections and distinctions between our method and prior works that also integrate ID (or Siamese encoder architecture) into MIM (or degraded input).
@@
Additional Results and Comparisons (Sec. E): Extended experiments comparing FSFM against other models, including ViT-based FAS and the base VLP model, CLIP.
@@ B Revealing the secrets of facial masking strategies in MIM
In the main paper, we explore various facial masking strategies for masked image modeling (MIM) in FSFM, with additional visualizations provided in Fig. 5, and validate the effectiveness of CRFR-P masking through ablation studies on downstream face security tasks. However, a critical question remains: how do different facial masking strategies affect the MIM pretrained model or its learned representations of real faces?
To address this, we quantitatively and qualitatively analyze the properties of attention maps. Given that most MIM pretrained models, including ours, are built on the Vision Transformer (ViT) architecture [28], whose main component, the attention mechanism [98], is naturally interpretable [115]. Here, we adopt the naive MAE (FSFM without the ID network) and follow its settings [45] with ViT-B/16 as the encoder and a 75% masking ratio. We conduct self-supervised pretraining on real face images from FF++_o [84] (the default dataset for our ablations). We alter only the masking strategy, encompassing simple random masking, Fasking-I, FRP, CRFR-R, and CRFR-P, and we examine the following aspects of attention heads in the pretrained models: 1) mean attention distance to measure the flow of local and global facial information; 2) KL divergence to investigate the diversity of attention; 3) visualized attention maps to uncover key regions of focus.
@@ B.1 Local or global?
To explore whether the pretrained model attends to faces locally or globally, we calculate the mean attention distance [28] in each attention head across all blocks/layers, as shown in Fig. 6 (top). The model (MAE ViT-B/16 encoder) pretrained with simple random masking tends to focus on local information in the lower blocks and shifts toward global attention in the deeper blocks, similar to the supervised model [28]. Fasking-I primarily aggregates global information as the visible patches predominantly consist of broad backgrounds and skin. FRP also causes large mean attention distances, but they are slightly smaller than those of Fasking-I, mainly because visible patches in FRP are more evenly distributed across all facial regions. CRFR-R fully masks a facial region before applying random masking, which encourages attention to different regions. Consequently, CRFR-R has more global attention in the middle (3rd to 8th) blocks than the simple random masking counterparts. Compared with CRFR-R, CRFR-P masks the remaining regions proportionally instead of randomly, making the 1st block more global w.r.t. the more unmasked regions. Compared with FRP, CRFR-P fully masks a region before applying proportional masking, exposing more patches within other regions at the same masking ratio, thus leading to more local attention than FRP.
Overall, the model pretrained with CRFR-P exhibits well-distributed mean attention distances across all blocks, indicating a synergistic effect of FRP and CRFR-R, enabling appropriate attention to both local and global information.
@@ B.3 Key regions?
To uncover which regions of real faces are critical to the pretrained model, we visualize the mean attention map of the last block and overlay it onto the input face in Fig. 7, as the pretext decoder or downstream head follows the last block. We can observe significant attention differences in the salient regions across different masking strategies. At first glance, the attention regions from simple random masking appear to cover the entire face. However, it predominantly highlights the skin, which can be easily recovered from visible neighboring patches, ignoring more challenging key regions. This suggests that the model solves face reconstruction through shortcuts instead of learning meaningful features. Similarly, the attention in Fasking-I is distributed across skin and background regions, as expected. While FRP and CRFR-R activate more attention areas, they still struggle to focus on key facial regions. In contrast, CRFR-P highlights attention across key regions like the eyes, eyebrows, and nose, indicating that the pretrained model tackles the challenge head-on: focusing on meaningful region-level features beyond just low-level pixels of real faces.
@@
| ViT-B [28]* | Sup.(IN) | 5.00 | 97.29 | 11.51 | 95.61 | 10.60 | 94.03 | 17.35 | 88.35 | 11.12 | 93.82 | |||||
| MAE [45] ViT-B* | SSL.(LN) | 8.17 | 97.37 | 27.91 | 77.90 | 18.02 | 91.54 | 25.36 | 80.76 | 19.86 | 86.89 | |||||
| DINO [11] ViT-B* | SSL.(LN) | 7.92 | 97.09 | 21.28 | 87.79 | 21.35 | 84.48 | 17.44 | 89.75 | 17.00 | 89.78 | |||||
| MCF [103] ViT-B* | SSL.(LFc) | 6.33 | 98.29 | 21.40 | 86.72 | 13.13 | 95.15 | 16.76 | 89.59 | 14.41 | 92.44 | |||||
| FSFM ViT-B (Ours)* | SSL.(VF2) | 6.58 | 97.43 | 4.30 | 99.10 | 14.63 | 90.96 | 10.02 | 96.23 | 8.88 | 95.93 | |||||
@@
| ViTranZFAS [33] (IJCB’21)* | Init.(IN) | 10.95 | 95.05 | 14.33 | 92.10 | 16.64 | 85.07 | 15.67 | 89.59 | 14.40 | 90.45 | |||||
| TransFAS [105] (TBIOM’22)* | Init.(IN) | ✓ | 7.08 | 96.69 | 9.81 | 96.13 | 10.12 | 95.53 | 15.52 | 91.10 | 10.63 | 94.86 | ||||
| TTN-S [106] (TIFS’22)* | Init.(IN) | ✓ | 9.58 | 95.79 | 9.81 | 95.07 | 14.15 | 94.06 | 12.64 | 94.20 | 11.55 | 94.78 | ||||
| DiVT-V [67] (WACV’23)* | Init.(IN) | ✓ | ✓ | 10.00 | 96.64 | 14.67 | 93.08 | 5.71 | 97.73 | 18.06 | 90.21 | 12.11 | 94.42 | |||
| TTDG-V [134] (CVPR’24)* | Init.(IN) | ✓ | 4.16 | 98.48 | 7.59 | 98.18 | 9.62 | 98.18 | 10.00 | 96.15 | 7.84 | 97.75 | ||||
@@
C Connection and analysis of ID in FSFM
C.1 Target view and network design
Based on our ablations (in the main paper) of downstream face security tasks, FSFM performs better when using full patches as the target view alongside Siamese rep decoders. By predicting the entire face representation from visible parts, the ID network aligns global and local views of the same face. In light of this, FSFM structures the encoded space with semantically complete facial representations through ”local-to-global” correspondences [11], which endows the encoder with strong facial discriminability.
C.2 Data augmentation
Most ID methods [14, 44, 16, 36, 15, 11] rely on strong data augmentations, including spatial and color transforms, to prevent trivial solutions. For MIM, applying strong augmentations, such as color enhancements, is suboptimal [45], as masking corruption itself effectively regularizes the pretext task. This is further evident in methods [119, 64, 55] that combine MIM and ID, where only simple augmentations—random size cropping or flipping—are applied to the masked input of the online branch, while strong or simple augmentations are used for the full (unmasked) target view.
@@ C.2 Data augmentation
C.3 Loss function for ID
In FSFM, we observe that NCS outperforms the contrastive loss InfoNCE, even though methods combining MIM and ID [64, 119, 55, 103] favor the latter. We speculate this is because, in large-scale pretraining on real faces, the inter-image contrast introduced by negative sample pairs—pushing one real face away from others—does not help our model learn representations that are beneficial for face security tasks. Thus, we adopt asymmetric NCS by default to learn intra-face correspondences by matching the online anchor view with the target view of the same sample.
D More implementation details
D.1 pretraining settings
D.2 Finetuning settings in downstream tasks
Deepfake Detection We use the c23 (HQ) version of FF++ [84] with official train/val splits for finetuning. We sample 128 frames per real video (the original YouTube subset) and 32 frames per forgery video (including Deepfakes, Face2Face, FaceSwap, and NeuralTextures subsets). We follow the official test split in other unseen datasets for testing, including CDFV2 [65], DFDC [27], DFDCp [26], and WDF [137]. We sample 32 frames per testing video. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224×224. As WDF already provides 224×224 facial images, we directly use its test set. We add only one linear layer as the binary classifier after averaging all non-CLS token features. We set the batch size to 64, the base learning rate to 2.5e-4, and the finetuning epochs to 10 (50 for ViT-B Scratch). Other settings adhere to the MAE ImageNet finetuning recipe [45].
Face Anti-Spoofing In the main paper, we adopt the 0-shot MCIO setting (Protocol 1) in [54] and include CelebA-Spoof [124] as supplementary data for FAS finetuning. We set the batch size to 12 for each training domain. We append the MLP head after averaging all non-CLS token features instead of using CLS ones [54], to align with other face security tasks. Additionally, for a fair comparison with other visual-only ViT-based methods, we additionally follow [134] and report the best performance without including the supplementary data, as presented in Sec. E.1.
Diffusion Face Forgery Detection For the training set, we sample 32 frames from each real video (the original YouTube subset) and each forgery video (the Deepfakes subset) from FF++ (c23) [84]. For validation and testing sets, we follow the splits provided by the DiFF benchmark [18]. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224×224. We add one linear layer as the binary classifier after averaging all non-CLS token features. We set the batch size to 256, the base learning rate to 5e-4, and the finetuning epochs to 50. Other settings adhere to the MAE ImageNet finetuning recipe [45].
E Additional experimental results
E.1 Comparison with ViT-based FAS
In a fair comparison with visual-only ViT-based FAS methods, our method also significantly outperforms all base models, as shown in Tab. 5. FSFM surpasses most counterparts and ranks second in average metrics. TTDG-V [134], which introduces test-time domain generalization and explicit optimization goals for FAS, performs better than ours in two out of four target domains (M and I). While optimizing for a specific downstream task is beyond the scope of this study, incorporating special auxiliary supervision or domain generalization (DG) techniques into our pretrained model may further improve its generalization ability for face presentation attack detection.
E.2 Comparison with CLIP

### v3

@@ 4.3 Face anti-spoofing
Comparison with Base Models Our FSFM achieves the best average performance, with several key observations: 1) Simple finetuning on ViT-B Sup(IN) enhances FAS generalization, as noted in [54, 135]. 2) General SSL methods, including MIM-based MAE [45] and ID-based DINO [11], perform worse than ViT-B Sup(IN), with MAE showing notable declines, echoing the claims in [121]. 3) With large-scale facial data for SSL, MCF outperforms MAE and DINO for FAS. However, similar to DfD, it still falls short of the ViT-B Sup(IN) baseline in average metrics, once again underscoring the gap between normal face analysis and security tasks. 4) Overall, our FSFM improves the generalizability of ViT for cross-domain FAS, effectively modeling credible features of live (real) faces.
Comparison with SOTA Specialized Methods Our FSFM ViT-B exhibits superior performance over domain generalization (DG) FAS methods using various CNNs. Notably, this comparison adheres to the baseline [54]: although supplementary data (CelebA-Spoof [125]) is included for finetuning ViT-B, no auxiliary supervision or DG techniques, e.g., depth map (DM) or adversarial learning (AL), are employed—only a standard cross-entropy loss. Additionally, we follow other ViT-based FAS methods [135, 67, 107] without including CelebA-Spoof, with comparisons provided in the supplementary material.
@@ 4.5 Ablation studies
In ablations, unless otherwise stated, we pretrained the ViT-B/16 on the FF++_o dataset, which includes 0.11M real face images from 720 training and 140 validation videos of FF++ (c23) YouTube subset [84]. We present the average generalization results on DfD and FAS in Tab. 4.
@@
Effect of Key Designs 1) Blocks of & A lightweight 2-layer Rep decoder in both the online and target branches (2&2) achieves a better balance between complexity and generalization, outperforming 1&1 and 3&3. Omitting the Rep decoder (0&0) or adding it only to the online branch (2&0) degrades performance, supporting the importance of bridging feature distribution in a disentangled space. 2) Data Augmentation FSFM performs best without data augmentation (none&none). Unlike many SSL methods, applying simple augmentation in the MIM network (crop+flip&none) or target view (none&crop+flip) reduces performance, indicating that CRFR-P masking introduces sufficient spatial regularization and that preserving target representations from original faces aids robustness. 3) Loss for ID Network Pretraining with asymmetric negative cosine similarity (NCS) i.e. Eq. 4 proves more effective than the BYOL-like [36] MSE and the widely used InfoNCE, as the latter pushes different real faces (negative samples) apart that may not be suitable for face security tasks.
@@ 4.6 Qualitative analysis
To illustrate the superiority of the facial features learned by FSFM, we visualize GradCAM [86] activation maps for detecting forgery and spoofing, compared to Sup(IN) and MAE baselines. 1) DfD In Fig. 4 (a), FSFM heatmaps more accurately identify forgery-relevant artifacts in FF++ [84], e.g., mouth-modified F2F and NT. 2) FAS In Fig. 4 (b), FSFM predominantly captures spoof-specific clues in the MCIO cross-domain testings: high-frequency paper textures (M-Paper), photo cut edges (C-Photo), inconsistent reflections between facial regions (I-Print), and moiré patterns in video (O-Replay). These visualizations show that FSFM effectively responds to anomalies diverging from the suggested 3C in real/live faces, shedding light on improving the generalization performance of face security tasks.
@@ 5 Conclusion
In this work, we propose a self-supervised pretraining framework to learn fundamental and general representations of real faces, FSFM. We leverage the synergy between masked image modeling and instance discrimination to empower both local and global perception, promoting intra-region consistency, inter-region coherency, and local-to-global correspondence. We show that our FSFM transfers better than previous pretrained models on several face security tasks, including cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion face forgery detection. Notably, as a vanilla ViT, our model surpasses SOTA task-specific methods on generalization performance. Finally, we hope that our work can contribute to the face security and analysis community.
Acknowledgements
This work was supported in part by the National Key R&D Program of China under Grant 2023YFB2904000 and 2023YFB2904001, in part by the National Natural Science Foundation of China under Grant 62032021, U2436206, 62372406, and 62372402, and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant LZ25F020005 and LD24F020010. The authors would like to thank all anonymous reviewers for their insightful comments on this paper. Besides, the first author wishes to thank his late grandmother, Juxian Huang, for her spiritual encouragement at the beginning of this work.
@@ A Overview
This supplementary material provides additional insights, details, and results to support our FSFM framework comprehensively, structured as follows:
Facial Masking Strategies in Masked Image Modeling (MIM) (Sec. B): We delve into the impact of different facial masking strategies on naive MAE, including quantitative and qualitative analysis of attention differences.
Instance Discrimination (ID) in FSFM (Sec. C): We highlight the connections and distinctions between our method and prior works that also integrate ID (or Siamese encoder architecture) into MIM (or degraded inputs).
@@
Additional Results and Comparisons (Sec. E): Extended experiments comparing FSFM against other models like ViT-based FAS and the base vision-language pertaining (VLP) model, CLIP.
@@ B Revealing the secrets of facial masking strategies in MIM
In the main paper, we explore various facial masking strategies for masked image modeling (MIM) in FSFM, with additional visualizations provided in Fig. 5, and validate the effectiveness of CRFR-P masking through ablation studies on downstream face security tasks. However, a critical question remains: how do different facial masking strategies affect the MIM-pretrained model or its learned representations of real faces?
To address this, we quantitatively and qualitatively analyze the properties of attention maps. Given that most MIM-pretrained models, including ours, are built on the Vision Transformer (ViT) architecture [28], where the main component, the attention mechanism [99], is naturally interpretable [116]. Here, we adopt the naive MAE (the MIM network in FSFM) and follow its settings [45] with ViT-B/16 as the encoder and a 75% masking ratio. We conduct self-supervised pretraining on real face images from FF++_o [84] (the default dataset for our ablations). We alter only the masking strategy, encompassing simple random masking, Fasking-I, FRP, CRFR-R, and CRFR-P, and examine the following aspects of attention heads in the pretrained models: 1) mean attention distance to measure the flow of local and global facial information; 2) Kullback-Leibler (KL) divergence to investigate the diversity of attention; 3) visualized attention maps to uncover key regions of focus.
@@ B.1 Local or global?
To explore whether the pretrained model attends to faces locally or globally, we calculate the mean attention distance [28] in each attention head across all blocks/layers, as shown in Fig. 6 (Top). The model (MAE ViT-B/16 encoder) pretrained with simple random masking tends to focus on local information in the lower blocks and shifts toward global attention in the deeper blocks, similar to the supervised model [28]. Fasking-I primarily aggregates global information as the visible patches predominantly consist of broad backgrounds and skin. FRP also causes large mean attention distances, but these are slightly smaller than those of Fasking-I, mainly because visible patches in FRP are more evenly distributed across all facial regions. CRFR-R fully masks a facial region before applying random masking, which encourages attention to different regions, consequently resulting in more global attention in the middle (3rd to 8th) blocks compared to the simple random masking counterparts. Compared with CRFR-R, CRFR-P masks the remaining regions proportionally instead of randomly, making the 1st block more global w.r.t. the more unmasked regions. Compared with FRP, CRFR-P fully masks a region before applying proportional masking, which exposes more patches within other regions at the same masking ratio, thus leading to more local attention than FRP.
Overall, the model pretrained with CRFR-P exhibits well-distributed attention distances across all blocks, indicating a synergistic effect of FRP and CRFR-R, enabling appropriate attention to both local and global information.
@@ B.3 Key regions?
To uncover which regions of real faces are critical to the pretrained model, we visualize the mean attention map of the last block and overlay it onto the input face in Fig. 7, as the pretext decoder or downstream head follows the last block. We can observe significant attention differences in salient regions across different masking strategies. At first glance, the attention regions from simple random masking appear to cover the entire face. However, it predominantly highlights the skin, which can be easily recovered from visible neighboring patches, while ignoring more challenging key regions. This suggests that the model solves face reconstruction through shortcuts instead of learning meaningful features. Similarly, the attention in Fasking-I is distributed across skin and background regions, as expected. While FRP and CRFR-R activate more attention areas, they still struggle to focus on key facial regions. In contrast, CRFR-P highlights attention across key regions like the eyes, eyebrows, and nose, indicating that the pretrained model tackles the challenge head-on: focusing on meaningful region-level features beyond just low-level pixels of real faces.
@@
C Connection and analysis of ID in FSFM
C.1 Target view and network design
Based on our ablations (in the main paper) of downstream face security tasks, FSFM performs better when using full patches as the target view alongside Siamese rep decoders. By predicting the entire face representation from visible parts, the ID network aligns global and local views of the same face. In light of this, FSFM structures the encoded space with semantically complete facial representations through “local-to-global” correspondences [11], which endows the encoder with strong facial discriminability.
C.2 Data augmentation
Most ID methods [14, 44, 16, 36, 15, 11] rely on strong data augmentations, including spatial and color transforms, to prevent trivial solutions. For MIM, applying strong augmentations such as color enhancements is suboptimal [45], as masking corruption itself effectively regularizes the pretext task. This is further evident in methods [120, 64, 55] that combine MIM and ID, where only simple augmentations—random size cropping or flipping—are applied to the masked input of the online branch, while strong or simple augmentations are used for the full (unmasked) target view.
C.3 Loss function for ID
In FSFM, we observe that NCS outperforms the contrastive loss InfoNCE, even though methods combining MIM and ID [64, 120, 55, 104] favor the latter. We speculate this is because, in large-scale pretraining on real faces, the inter-image contrast introduced by negative sample pairs——pushing one real face away from others—does not help our model learn representations beneficial for face security tasks. Thus, we adopt asymmetric NCS by default to learn intra-face correspondences by matching the online anchor view with the target view of the same sample.
| Base Model | Pretrain or Init | Intra-FF++ (c23) | Intra-FF++ (c40) | ||
| F-AUC | V-AUC | F-AUC | V-AUC | ||
| ViT-B [28] | Scratch | 70.06 | 72.98 | 70.36 | 74.33 |
| ViT-B [28] | Sup(IN) | 97.30 | 98.80 | 88.65 | 93.22 |
| MAEE [45] ViT-B | SSL(LN) | 98.64 | 99.61 | 91.35 | 94.46 |
| DINO [11] ViT-B | SSL(LN) | 98.35 | 99.39 | 89.93 | 94.31 |
| MCF [104] ViT-B | SSL(LFc) | 97.84 | 99.19 | 89.81 | 94.20 |
| FSFM ViT-B (Ours) | SSL(VF2) | 97.74 | 99.03 | 92.08 | 95.41 |
@@ C.3 Loss function for ID
| ViT-B [28]* | Sup(IN) | 5.00 | 97.29 | 11.51 | 95.61 | 10.60 | 94.03 | 17.35 | 88.35 | 11.12 | 93.82 | |||||
| MAE [45] ViT-B* | SSL(LN) | 8.17 | 97.37 | 27.91 | 77.90 | 18.02 | 91.54 | 25.36 | 80.76 | 19.86 | 86.89 | |||||
| DINO [11] ViT-B* | SSL(LN) | 7.92 | 97.09 | 21.28 | 87.79 | 21.35 | 84.48 | 17.44 | 89.75 | 17.00 | 89.78 | |||||
| MCF [104] ViT-B* | SSL(LFc) | 6.33 | 98.29 | 21.40 | 86.72 | 13.13 | 95.15 | 16.76 | 89.59 | 14.41 | 92.44 | |||||
| FSFM ViT-B (Ours)* | SSL(VF2) | 6.58 | 97.43 | 4.30 | 99.10 | 14.63 | 90.96 | 10.02 | 96.23 | 8.88 | 95.93 | |||||
@@
| ViTranZFAS [33] (IJCB’21)* | Init(IN) | 10.95 | 95.05 | 14.33 | 92.10 | 16.64 | 85.07 | 15.67 | 89.59 | 14.40 | 90.45 | |||||
| TransFAS [106] (TBIOM’22)* | Init(IN) | ✓ | 7.08 | 96.69 | 9.81 | 96.13 | 10.12 | 95.53 | 15.52 | 91.10 | 10.63 | 94.86 | ||||
| TTN-S [107] (TIFS’22)* | Init(IN) | ✓ | 9.58 | 95.79 | 9.81 | 95.07 | 14.15 | 94.06 | 12.64 | 94.20 | 11.55 | 94.78 | ||||
| DiVT-V [67] (WACV’23)* | Init(IN) | ✓ | ✓ | 10.00 | 96.64 | 14.67 | 93.08 | 5.71 | 97.73 | 18.06 | 90.21 | 12.11 | 94.42 | |||
| TTDG-V [135] (CVPR’24)* | Init(IN) | ✓ | 4.16 | 98.48 | 7.59 | 98.18 | 9.62 | 98.18 | 10.00 | 96.15 | 7.84 | 97.75 | ||||
@@
D More implementation details
D.1 Pretraining settings
D.2 Finetuning settings in downstream tasks
Deepfake Detection We use the c23 (HQ) version of FF++ [84] with official train/val splits for finetuning. We sample 128 frames per real video (the original YouTube subset) and 32 frames per forgery video (including Deepfakes, Face2Face, FaceSwap, and NeuralTextures subsets). We follow the official test split in other unseen datasets for testing, including CDFV2 [65], DFDC [27], DFDCp [26], and WDF [138]. We sample 32 frames per testing video. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224224. As WDF already provides 224224 facial images, we directly use its test set without processing. We add only one linear layer as the binary classifier after averaging all non-[CLS] token features. We set the batch size to 64, the base learning rate to 2.5e-4, and the finetuning epochs to 10 (50 for ViT-B Scratch). Other settings adhere to the MAE ImageNet finetuning recipe [45].
@@ D.2 Finetuning settings in downstream tasks
Face Anti-Spoofing In the main paper, we adopt the 0-shot MCIO setting (Protocol 1) in [54] and include CelebA-Spoof [125] as supplementary data for FAS finetuning. We set the batch size to 12 for each training domain. We append the MLP head after averaging all non-[CLS] token features instead of using [CLS] ones [54], to align with other face security tasks. Additionally, for a fair comparison with other visual-only ViT-based methods, we additionally follow [135] and report the best performance without including the supplementary data, as presented in Sec. E.2.
Diffusion Face Forgery Detection For the training set, we sample 32 frames from each real video (the original YouTube subset) and each forgery video (the Deepfakes subset) from FF++ (c23) [84]. For validation and testing sets, we follow the splits provided by the DiFF benchmark [18]. We use DLIB [59] to extract faces (without alignment and parsing) and resize them to 224224. We add one linear layer as the binary classifier after averaging all non-[CLS] token features. We set the batch size to 256, the base learning rate to 5e-4, and the finetuning epochs to 50. Other settings adhere to the MAE ImageNet finetuning recipe [45].
E Additional experimental results
E.1 Comparison of intra-dataset DfD on FF++
While our primary objective focuses on cross-domain generalization for real-world applicability, we provide an intra-dataset evaluation of deepfake detection (DfD) on FaceForensics++ (FF++) [84], as presented in Tab. 5. The metrics on the FF++_c23 show that FSFM maintains comparable intra-set performance while significantly improving cross-dataset generalization. Moreover, when evaluated on the more challenging high-compression (c40) version, FSFM outperforms all baseline vision models, further demonstrating its robustness.
E.2 Comparison with ViT-based FAS
In a fair comparison with visual-only ViT-based face anti-spoofing (FAS) methods, our method also significantly outperforms all base models, as shown in Tab. 6. FSFM surpasses most counterparts and ranks second in average metrics. TTDG-V [135], which introduces test-time domain generalization and explicit optimization goals for FAS, performs better than ours in two out of four target domains (OCI→M and OCM→I). While optimizing for a specific downstream task is beyond the scope of this study, incorporating special auxiliary supervision or domain generalization (DG) techniques into our pretrained model may further improve its generalization ability for face presentation attack detection.
E.3 Comparison with CLIP